# Benchmarking Large Language Models for Cyberbullying Detection in Real-World YouTube Comments

### v1

@@ VIII Limitations
    Zero-shot prompts. All models were evaluated using identical zero-shot prompts without model-specific fine-tuning or prompt optimization. This setup prioritizes comparability but may underrepresent each model’s best-case performance.
@@
    Comment-level context. Each comment was moderated in isolation, without access to prior messages or conversation history. This limits the ability to detect context-dependent abuse, sarcasm, or coordinated behavior across threads.
@@
    Cultural bias in labeling. Human annotations were provided by two reviewers with similar linguistic and cultural backgrounds, which may have introduced bias in interpreting sarcasm, slang, or ambiguous phrasing.
@@
    Limited language scope. Although the dataset includes occasional comments in Arabic or Indonesian, it was primarily designed for English-language analysis. As such, results should not be generalized to multilingual performance.
@@
This study uses only publicly available YouTube comments. All user identifiers and any personally identifying details were removed or replaced with neutral placeholders before analysis, so no individual can be re-identified. Because the data are de-identified and pose minimal risk, no formal IRB review was required.
@@
The full, de-identified comment corpus, together with the human labels and model predictions, is openly available at https://github.com/Ammce/papers/tree/main/llm-cyberbullying-moderation%20. To minimise re-identification risk, user names, links, and YouTube comment IDs have been removed. The list of source‐video IDs can be shared with editors or qualified researchers under a non-disclosure agreement. Exact model prompts and API parameters used in this study are provided in the same repository under prompts/.

### v2

@@ VIII Limitations
    Zero-shot prompts: All models were evaluated using identical zero-shot prompts without any model-specific fine-tuning or prompt optimization. This setup maximizes comparability but may understate each model’s peak performance.
@@
    Comment-level context: Each comment was assessed in isolation, with no access to preceding messages or thread history. As a result, context-dependent abuse, sarcasm, or coordinated behavior across comments may go undetected.
@@
    Cultural bias in labeling: Human annotations came from two reviewers sharing similar linguistic and cultural backgrounds, which could bias judgments of sarcasm, slang, or ambiguous phrasing.
@@
    Limited language scope: Although a few comments were in Arabic or Indonesian, the dataset is overwhelmingly English. Therefore, our findings should not be generalized to truly multilingual moderation performance.
  5. 5.
    No BERT-based baseline: Traditional transformer classifiers like BERT and RoBERTa have already been widely studied in toxic language detection tasks [17, 18]. This work focuses instead on production-grade LLMs that are currently deployed or considered for content moderation at scale. Including legacy baselines would provide limited additional insight relative to the study’s primary goal of evaluating state-of-the-art zero-shot LLM performance.
@@
This study uses only publicly available YouTube comments. All user identifiers and any personally identifying details were removed or replaced with neutral placeholders before analysis, so no individual can be re-identified. Because the data are de-identified and pose minimal risk, no formal Institutional Review Board (IRB) review was required.
@@
The full, de-identified comment corpus, together with the human labels and model predictions, is openly available at https://github.com/Ammce/papers/tree/main/llm-cyberbullying-moderation%20. To minimize re-identification risk, user names, links, and YouTube comment IDs have been removed. The list of source‐video IDs can be shared with editors or qualified researchers under a non-disclosure agreement. Exact model prompts and API parameters used in this study are provided in the same repository under prompts/.